# {tocentry} Vibrational Energy Dissipation in Non-Contact Single-Molecule Junctions Governed by Local Geometry and Electronic Structure

### v1

@@ Discussion
Perspective and potential experimental validation. Finally, equilibrium correlation analysis and kinetic energy decay yield lifetimes in good agreement, which demonstrates the validity of the former. Equilibrium correlation analysis can yield important information, such as coupling densities and phononic spectral functions. Our method for disentangling electron-phonon and phonon-phonon contributions to the vibrational dynamics of adsorbed molecules enables a better understanding of electronic and vibrational mass effects, as well as effects due to the local bonding environment. It can improve the interpretability of vibrational line widths and energy dissipation observed in AFM, IETS, Raman, or infrared spectroscopy. If the relaxation of a vibrational mode is primarily governed by electron-phonon coupling, adjusting electronic properties, such as the tip material, functionalization, or the bias voltage, can enhance signal detection and reveal details about the coupling. If phonon-phonon coupling dominates, tuning temperature or using a different isotope (14C or 13C instead of 12C in CO, for instance) can shift vibrational frequencies and affect the coupling. In scanning probe experiments and single-molecule manipulation, where vibrational excitation can induce molecular motion or reactions, understanding electron-phonon and phonon-phonon relaxation supports design conditions that favor specific molecular transformations. Similarly, in catalysis, targeting specific vibrational modes can selectively trigger reactions, enhancing the control and efficiency of chemical processes.
@@ 3 Conclusion
The vibrational dynamics of CO-decorated copper tips that range from atomically flat to maximally sharp exhibit pronounced vibrational mode specificity, driven by the interplay of electron-phonon and phonon-phonon coupling, as well as significant geometric effects. Lifetimes due to electron-phonon coupling differ by two orders of magnitude between modes due to vibrational energies, the symmetry of the mode, and the DOS. Overall, CO vibrations of flatter tip geometries exhibit stronger electron-phonon relaxation, due to the increased energetic delocalization of the local electronic DOS. However, the high-energy IS mode does not show this trend due to the large energy window to excite electrons. Phonon-phonon lifetimes also decrease as the surface becomes flatter, although the IS mode does not follow this trend. The FT, FR, and SA modes are coupled to the substrate vibrations, leading to significant geometry dependence, while the IS mode is decoupled from the substrate vibrations. The sum of individual electron–phonon and phonon–phonon relaxation rates obtained from dynamic simulations tends to underestimate the total relaxation rate computed by the same method. This weak non-additivity arises from electron-phonon-mediated phonon-phonon coupling, which enhances phonon-phonon interactions. Finally, the lifetimes obtained from equilibrium correlation analysis and kinetic energy decay are in good agreement and also agree well with experiment.
Future work will explore the role of non-diagonal electronic friction on the mode selectivity and surface dependence. Incorporating a fully configuration-dependent electronic friction tensor will enable a more accurate investigation of non-additivity.46 Finally, extending the current framework to account for zero-point energy contributions will further improve the fidelity of atomic-scale friction models.
@@ 4 Methods
4.1 Modelling the PES of CO on Cu(111)
Due to the many-electron self-interaction error, the energies of the states responsible for bonding the CO molecule to the Cu surface are misrepresented. This results in an incorrect preference for the fcc-hollow site over the top site, which experimental studies have shown to be the correct adsorption site.47 Generalized Gradient Approximation (GGA) functionals fail to accurately describe CO adsorption on Cu and other metal surfaces.24, 25, 26 The Random Phase Approximation (RPA) has been proposed as a method to yield qualitatively correct adsorption sites, while the PBE0 functional shows a slight preference for the correct adsorption site.48 However, RPA calculations are unfeasible due to their high computational cost, while PBE0 is not ideal for modelling bulk metals since it tends to overestimate the delocalization of electrons, which can lead to an inaccurate description of metallic systems, particularly their electronic structure and density of states near the Fermi level.49, 50 In our work, we use the HSE06 screened hybrid functional,27 which treats short-range interactions with exact exchange and reduces the amount of exact exchange at long range. This helps mitigate some of the issues, such as over-delocalization of electrons and poor convergence in calculations.51 We combine this function with the MBD-NL long-range dispersion correction28—a combination that was previously shown to provide accurate adsorption structure and energetics.52 With this combination, we predict an accurate PES (see Table S1 in the Supporting Information). Calculations are performed using the FHI-aims quantum chemistry code53 with periodic boundary conditions and a repeated slab approach, including a dipole correction. The sharp and blunt tip geometries are placed on a four-layer slab for better convergence. Convergence tests and further details are shown in Section S1 of the Supporting Information.
@@ 4.2 Interatomic potential
We use the MACE equivariant neural network potential32 and train it at the HSE06+MBD-NL level of theory, which accurately describes the potential energy surface of CO on Cu(111) as well as CO adsorbed on an adatom. Using a training set of 621 data points, the potential achieves an energy error of /atom and a force error of , which was determined using k-fold cross-validation, where we split our dataset into five equally sized segments. The training data contains structures of all four surface geometries (sharp tip, blunt tip, slab adatom, slab). Details regarding hyperparameters, training/test splits, and validation during dynamics simulations are shown in Section S2 in the Supporting Information.
4.3 Vibration calculatons
@@ 4.4 Description of electron-phonon coupling
In first-order perturbation theory, we can express the elements of the electronic friction tensor associated with atomic degrees of freedom and as follows:35, 21
The electronic friction tensor associated with energy of a given vibration mode is calculated by summing over all possible excitations between effective single particle Kohn–Sham states and . The indices and indicate the initial and final state, respectively, and is the associated wave vector. The terms in the sum are given by the nonadiabatic coupling matrix elements (first row of Equation 7), the occupation factors and of the initial and final states (second row), and the conservation of energy (third row). The term accounts for the perturbing frequency of the vibration mode of interest. We calculate electronic friction with the FHI-aims quantum chemistry code.35 Convergence tests for electronic friction calculations can be found in Section 1.6 in the Supporting Information. We explicitly calculate electronic friction for C and O atoms. For the Cu atoms, we use isotropic electronic friction and a diagonal friction tensor. This friction value is calculated for a single atom in bulk copper (see Section S1.7 in the Supporting Information). Electron-phonon coupling calculations are done using the PBE54 exchange-correlation functional, as has been done in previous simulations of CO on Cu substrates.21, 35 Details can be found in Section S1.7 of the Supporting Information. We note that employing the HSE06 functional for electronic friction calculations in systems of this size is currently infeasible given available high-performance computing resources. Vibrational relaxation rates are determined from the friction tensor and the wass-weighted vibrational eigenmodes, , at the relevant vibrational frequency :
@@ 4.5 Molecular dynamics
We perform molecular dynamics (MD) simulations using the atomic simulation environment (ASE).55 A time step of is employed. Since we are interested in the dynamics of a single CO molecule, we conduct MD simulations in an extended unit cell containing a single adsorbed CO molecule, shown in Figure 1. The extended unit cell facilitates the back-folding of substrate phonon bands, enabling effective sampling of the metal phonon band structure.
To account for electron-phonon coupling, we use molecular dynamics with electronic friction, where the dynamics are governed by the Langevin equation:56
Here, represents the potential energy surface, which depends on the atomic positions . The term denotes the electronic friction tensor, while is the random force responsible for maintaining detailed balance at a given electronic temperature , as dictated by the second fluctuation-dissipation theorem.57 For the dynamics, the electronic friction tensor is calculated in the quasi-static limit with a perturbation frequency of zero. A comparison of lifetimes with and without a perturbing frequency can be found in Section S3.1 of the Supporting Information. During the calculations, the constant friction approximation is used for the electronic friction tensor, where the friction tensor values have been determined at the relaxed equilibrium geometry of the molecule on the surface.
To determine relaxation rates and lifetimes from the MD and MDEF trajectories, we use three different methods:
@@
For each system, we first perform equilibration at , followed by computing a single, long MD or MDEF trajectory (see Section S3.2 of the Supporting Information). To analyse the velocity time series, we use normal mode decomposition and determine the cross-spectrum using the Welch method. Thereby, the velocity time series is cut into multiple, partially overlapping segments. A window function—a Hann window in our case—is then applied to each segment before the cross-spectrum is calculated. The cross-spectra of the different segments are then averaged. This improves the smoothness of the cross-spectrum. The height and width of the peak(s) in the cross-spectrum are then determined by fitting a Lorentzian function (see Section 4 of the Supporting Information).
@@
For each system and vibrational mode, relaxation rates and lifetimes are extracted from ensemble averages over 64 MD or MDEF trajectories. We first equilibrate the system at . Then, we excite the vibrational mode of interest by a displacement amplitude equivalent to along the respective eigenvector. Using normal mode decomposition, we project the velocity time series onto the basis of the vibration modes. We subsequently monitor the temporal decay of the envelope of the ensemble-averaged kinetic energy. Assuming exponential decay—as would be the case for a damped harmonic oscillator—we determine relaxation rates via an exponential fit.
@@
Total relaxation rates are calculated by summing the individual relaxation rates associated with phonon-phonon and electron-phonon coupling. Phonon-phonon relaxation rates are derived from the inverse values listed in Table 2, which were determined through equilibrium correlation analysis. Electron-phonon relaxation rates are obtained from the inverse values in Table 3, which were directly calculated from the electronic friction tensor.

### v2

@@ Discussion
Finally, our work can improve the interpretability of vibrational line widths and energy dissipation observed in AFM, IETS, TERS, or IR spectroscopy. The spectral function calculated through equilibrium correlation analysis provides mode-specific linewidths and lifetimes, while the separation of electron–phonon and phonon–phonon contributions enables clearer interpretation of lineshapes and mode-specific enhancements, particularly in IR and TERS. By disentangling dissipation pathways, our method can facilitate cross-correlation between IR and IETS signals and help assign adsorption sites based on the adsorption environment. Understanding which relaxation pathway dominates can enable experimental tuning. If the relaxation of a vibrational mode is primarily governed by EPC, adjusting electronic properties, such as the tip material, functionalization, or the bias voltage, can enhance signal detection and reveal details about the coupling. If PPC dominates, tuning temperature or using a different isotope (14C or 13C instead of 12C in CO, for instance) can shift vibrational frequencies and affect the coupling. In scanning probe experiments and single-molecule manipulation, where vibrational excitation can induce molecular motion or reactions, understanding electron-phonon and phonon-phonon relaxation supports design conditions that favor specific molecular transformations. Similarly, in catalysis, targeting specific vibrational modes can selectively trigger reactions, enhancing the control and efficiency of chemical processes.
@@ 3 Conclusion
The vibrational dynamics of CO-decorated copper tips that range from atomically flat to maximally sharp exhibit pronounced vibrational mode specificity, driven by the interplay of EPC and PPC, as well as significant geometric effects. Lifetimes due to EPC differ by two orders of magnitude between modes due to vibrational energies, the symmetry of the mode, and the DOS. Overall, CO vibrations of flatter tip geometries exhibit stronger electron-phonon relaxation, due to the increased energetic delocalization of the local electronic DOS. However, the high-energy IS mode does not show this trend due to the large energy window to excite electrons. Phonon-phonon lifetimes also decrease as the surface becomes flatter, although the IS mode does not follow this trend. The FT, FR, and SA modes are coupled to the substrate vibrations, leading to significant geometry dependence, while the IS mode is decoupled from the substrate vibrations. The sum of individual electron–phonon and phonon–phonon relaxation rates tends to underestimate the total relaxation rate computed by the same method. This weak non-additivity arises from electron-phonon-mediated PPC, which enhances phonon-phonon interactions. Finally, the lifetimes obtained from equilibrium correlation analysis and kinetic energy decay are in good agreement and also agree well with experiment.
Future work will explore the role of non-diagonal electronic friction on the mode selectivity and surface dependence. Incorporating a fully configuration-dependent electronic friction tensor will enable a more accurate investigation of non-additivity.56 Finally, extending the current framework to account for zero-point energy contributions will further improve the fidelity of atomic-scale friction models.
@@ 4 Methods
4.1 Modelling CO on different adsorption environments
Electronic structure calculations are performed using DFT with the FHI-aims code,57 employing periodic boundary conditions and a repeated-slab approach including a dipole correction. The sharp and blunt tip geometries are placed on a four-layer slab for better convergence. Convergence tests and further details are shown in Section S1 of the Supporting Information.
Commonly used methods for first-principles surface simulations, such as Generalized Gradient Approximation (GGA) functionals, fail to accurately describe CO adsorption on Cu and other metal surfaces.26, 27, 28 Due to the many-electron self-interaction error, the energies of the states responsible for bonding the CO molecule to the Cu surface are misrepresented. This results in an incorrect preference for the fcc-hollow site over the top site, which experimental studies have shown to be the correct adsorption site.58 The Random Phase Approximation (RPA) has been proposed as a method to yield qualitatively correct adsorption sites, while the PBE0 functional shows a slight preference for the correct adsorption site.59 However, RPA calculations are unfeasible due to their high computational cost, while PBE0 is not ideal for modelling bulk metals since it tends to overestimate the delocalization of electrons, which can lead to an inaccurate description of metallic systems, particularly their electronic structure and density of states near the Fermi level.60, 61
In our work, we use the HSE06 screened hybrid functional,29 which treats short-range interactions with exact exchange and reduces the amount of exact exchange at long range. This helps mitigate some of the issues, such as over-delocalization of electrons and poor convergence in calculations.62 We combine this function with the MBD-NL long-range dispersion correction30—a combination that was previously shown to provide accurate adsorption structure and energetics.63 With this combination, we predict an accurate PES (see Table S1 in the Supporting Information). Geometry optimizations were performed for all systems with the four bottom layers frozen, prior to determining other properties (PDOS, EPC, PPC, etc.).
@@ 4.2 Interatomic potential
We use the MACE equivariant neural network potential34 and train it at the HSE06+MBD-NL level of theory, which accurately describes the potential energy surface of CO on Cu(111) as well as CO adsorbed on an adatom. Using a training set of 621 data points, the potential achieves an energy error of /atom and a force error of , which was determined using k-fold cross-validation, where we split our dataset into five equally sized segments. The training data contains structures of all four adsorption environments (sharp tip, blunt tip, slab adatom, slab). Details regarding hyperparameters, training/test splits, and validation during dynamics simulations are shown in Section S2 in the Supporting Information.
4.3 Vibration calculations
@@ 4.4 Description of electron-phonon coupling
In first-order perturbation theory, we can express the (33) atomwise block of the electronic friction tensor associated with atomic degrees of freedom and as follows:38, 23
The electronic friction tensor associated with energy of a given vibration mode is calculated by summing over all possible excitations between effective single particle Kohn–Sham states and . The indices and indicate the initial and final electronic state, respectively, and is the associated wave vector. The terms in the sum are given by the nonadiabatic coupling matrix elements (first row of Equation 7), the occupation factors and of the initial and final states (second row), and the conservation of energy (third row). Since we are dealing with electrons, the occupation factors follow the Fermi–Dirac distribution, which introduces the temperature dependence. The term accounts for the perturbing frequency of the vibration mode of interest. We calculate electronic friction with the electronic structure code FHI-aims.38 Convergence tests for electronic friction calculations can be found in Section 1.6 in the Supporting Information. We explicitly calculate electronic friction for C and O atoms. For the Cu atoms, we use isotropic electronic friction and a diagonal friction tensor. This friction value is calculated for a single atom in bulk copper (see Section S1.7 in the Supporting Information). EPC calculations are performed using the PBE64 exchange-correlation functional, as has been done in previous simulations of CO on Cu substrates.23, 38 Details can be found in Section S1.7 of the Supporting Information. We note that employing the HSE06 functional for electronic friction calculations in systems of this size is currently infeasible given available high-performance computing resources. Vibrational relaxation rates are determined from the friction tensor and the mass-weighted vibrational eigenmodes, , at the relevant vibrational frequency :
@@ 4.5 Molecular dynamics
We perform MD simulations using the atomic simulation environment (ASE).65 A time step of is employed. Since we are interested in the dynamics of a single CO molecule, we conduct MD simulations in an extended unit cell containing a single adsorbed CO molecule, shown in Figure 1. The extended unit cell facilitates the back-folding of substrate phonon bands, enabling effective sampling of the phonon band structure.
To account for EPC, we use MDEF, where the dynamics are governed by the Langevin equation:66
Here, represents the potential energy surface, which depends on the atomic positions . The term denotes a (33) atomwise block of the electronic friction tensor, while is the random force vector responsible for maintaining detailed balance at a given electronic temperature , as dictated by the second fluctuation-dissipation theorem.67 For the dynamics, the electronic friction tensor is evaluated in the harmonic approximation at zero perturbation frequency, based on a fixed nuclear geometry. A comparison of lifetimes with and without a perturbing frequency can be found in Section S3.1 of the Supporting Information. During the calculations, the constant friction approximation is used for the electronic friction tensor, where the friction tensor values have been determined at the relaxed equilibrium geometry of the molecule on the surface. To determine relaxation rates and lifetimes from the MD and MDEF trajectories, we use three different methods:
@@
For each system, we first perform equilibration at , followed by computing a single, long MD or MDEF trajectory. To analyse the velocity time series, we use normal mode decomposition and determine the cross-spectrum using the Welch method. Thereby, the velocity time series is cut into multiple, partially overlapping segments. A window function—a Hann window in our case—is then applied to each segment before the cross-spectrum is calculated. The cross-spectra of the different segments are then averaged. This improves the smoothness of the cross-spectrum. The height and width of the peak(s) in the cross-spectrum are then determined by fitting a Lorentzian function. A detailed workflow is provided in Section S3.2 of the Supporting Information.
@@
For each system and vibrational mode, relaxation rates and lifetimes are extracted from ensemble averages over 64 MD or MDEF trajectories. We first equilibrate the system at . Then, we excite the vibrational mode of interest by a displacement amplitude equivalent to along the respective eigenvector. Using normal mode decomposition, we project the velocity time series onto the basis of the vibration modes. We subsequently monitor the temporal decay of the envelope of the ensemble-averaged kinetic energy. Assuming exponential decay—as would be the case for a damped harmonic oscillator—we determine relaxation rates via an exponential fit. See Section S3.3 of the Supporting Information for full details.
@@
Total relaxation rates are calculated by summing the individual relaxation rates associated with phonon-phonon and EPC. Phonon-phonon relaxation rates are derived from the inverse values listed in Table 2, which were determined through equilibrium correlation analysis. Electron-phonon relaxation rates are obtained from the inverse values in Table 3, which were directly calculated from the electronic friction tensor. See Section S3.4 of the Supporting Information for the computational workflow.
@@ 6 Data availability
7 Supporting information
Extended description of electronic structure calculation settings and convergence tests; details of interatomic potential parametrization; technical information on lifetime calculations and tabulated numerical data; comparison of results with other coinage metals; additional analysis of molecular dynamics trajectories, including extended results on electron–phonon and phonon–phonon spectral functions.